# A Network Science Approach to Granular Time Series Segmentation

### v1

@@ Abstract
keywords:
time series , time series segmentation , graph neural networks , graphs , transformations
journal: Computer Communications
[…]
[inst1]organization=Jozef Stefan Institute, Department of Communication Technologies,addressline=Jamova 39, postcode=1000, city=Ljubljana, country=Slovenia
@@ 1 Introduction
In today’s digital era, time series (TS) are one of the most frequently encountered data types. Low-cost, high-resolution sensors have become a key part in mobile devices, industrial monitoring and healthcare, producing large amounts of TS data that capture the temporal patterns and essential statistical properties of various processes or phenomena [1]. With the rapid progress in automatic data acquisition and storage, we are now faced with the challenge of data that far exceeds our comprehension. TS analysis methods such as clustering, classification or regression, are therefore essential for effective and efficient processing of TS data [2, 3, 4, 5]. Although these methods are crucial, they often benefit from first segmenting the TS into smaller, more manageable parts, through a process known as Time Series Segmentation (TSS). It first involves identifying critical time points dividing the TS into smaller segments and second label them based on activity. This segmentation step can significantly simplify subsequent tasks such as classification or anomaly detection and provide more targeted insights. For example, in the field of telecommunications [6] the increasing volume of the network traffic data is a constant challenge. With TSS, we can segment network traffic, making it easier to detect patterns and respond swiftly to potential issues, such as peak usage periods, while optimizing bandwidth efficiency.
@@
In recent years, deep learning (DL) [7, 8] architectures have been introduced as a TSS method, however they have not yet reached the same level of performance in TSS as in other fields such as computer vision. Many of these approaches rely on sliding-window techniques, where the TS is partitioned into evenly sized segments that are individually labeled before being recombined. While sliding-window methods yield satisfactory results, they limit the granularity of segmentation due to fixed window sizes and strides.
@@
A different, less mainstream approach to TS analysis, explores graphs as a viable TS representation [9, 10, 11, 12] where each value in the raw TS is represented as a node. Transforming TS into graphs not only preserves the temporal order of the data, but also reveal interactions and dependencies between non consecutive points, which is important for tasks such as TSS. This transformations enables the adoption of Graph Neural Networks (GNNs), which are particularly well-suited to learn complex structural features. GNNs can help identify the critical transition points that distinguish different segments, and are able to capture subtle temporal dependencies and variations that often precede significant changes in the TS. Such representations based on node density and message passing between node neighborhoods naturally eliminate the fixed window sizes and strides from DL.
@@
By considering the combination of TS graph representations and GNNs, and taking inspiration from the dense labeling approach in computer vision where every pixel is assigned a class label [13], we can reformulate the TSS challenge as a node-level classification problem. In our graph representation of TS data, dense labeling entails assigning a specific label to each node, ensuring that even the most subtle transitions are detected and appropriately classified. In this work we aim to conduct a comprehensive and robust exploration of the different graph representations for TSS on the TSSB benchmark datasets in combination with GNNs. Based on the findings of this exploration, we propose a new approach that combines the Weighted Dual Perspective Visibility Graph (WDPVG) transformation with a GAT-based GNN architecture. The contributions of this paper are as follows:
  • 1.
@@ 2 Related work
Time series segmentation is an area of research in the TS analysis field, which received less attention compared to the other related TS tasks. However, its importance is emphasized by its application in various domains, such as healthcare, finance, energy, traffic and networking. We can divide it into three main categories depending on its application: Change Point Detection (CPD)[14][15][16], Windowed Segment Classification (WSC)[7] and Dense Labeling or Precise Segmentation (DLPS)[17]. An example of the three approaches can be found in Figure 1, where Figure 1a) shows the ground truth TS example with 4 different segments, Figure 1b) shows the output of the CPD method, Figure 1c) shows the output of the WSC segmentation and the last Figure 1d) shows the DLPS segmentation.
@@
Early research predominantly concentrated on identifying boundaries where abrupt changes occur, known as Change Points (CPs). Arif et al. [15] analyzed 25 years of Malaysian climate data to identify significant shifts in climate patterns. Their methodology employed two statistical techniques: Cumulative SUM (CUSUM) to detect shifts and bootstrapping to confirm the significance of those shifts. In contrast, Ermshaus et al.[14] treated CPD primarily as a preprocessing step for TSS. Their hierarchical approach recursively partitions a univariate time series into two segments. For each potential split point, a binary k-NN classifier is trained, and the split point yielding the highest accuracy is selected as the change point. ClaSP[14] further automates this process by determining hyperparameters from the data through self-supervision and specialized algorithms, thus removing the need for manual parameter tuning. Because traditional CPD methods often struggle with subtle changes in autocorrelation structures and may exhibit high false-alarm rates, De Ryck et al.[18] addressed these challenges by employing a deep learning autoencoder. Their approach learns a time-invariant representation of the time series, focusing on stable features across the data. To detect change points, they calculate a dissimilarity measure based on Euclidean distances between consecutive windows’ feature representations. Notable increases in this measure indicate potential change points.
@@
However, a key limitation of CPD lies in its inability to classify the underlying nature of each segment. Consequently, TSS has evolved into a classification problem to overcome this gap. The most common strategy is WSC, in which the time series is split into fixed-length windows, each labeled independently. For instance, Perslev et al.[7] employed a CNN-based U-Time architecture to segment sleep stages in EEG signals. A drawback, however, is that segmentation granularity depends on the chosen window size. Reducing the window to a single time stamp yields the most granular approach, referred to as DLPS. Gaugel et al.[17] introduced a deep learning model that merges sliding window and dense labeling techniques by employing CNNs for intra-window feature extraction, LSTM networks for inter-window contextual understanding, and a refinement module for dense labeling. Evaluated on multivariate time series with their model achieved a segmentation accuracy of approximately 96%.
Compared to the presented related work, despite notable similarities, majority of the works focus on the CPD, while DLPS is mostly done on multivariate TS. To the best of our knowledge, this work is the first formulating the segmentation of univariate TS data as DLPS, through the graph dimensionality expansion representation of TS and GNNs.
@@ 3 Problem formulation
TSS as a fundamental task in time series analysis involves partitioning a sequence into semantically meaningful segments. In this work, we formulate TSS as a point-wise classification problem, also referred to as DLPS, where each time step in a time series is assigned a categorical label indicating its segment membership. This formulation aligns well with graph node classification tasks [19]. By representing the TS as a graph, where each time point corresponds to a graph node, the task of segmenting the series translates to assigning labels to these nodes.
@@ 4 Time series to graph transformations
In this section, we describe the main data preparation step, how to transform TS into graphs. We define the transformation $T()$ that transforms the input TS $S$ to its corresponding graph $G$: $G=T(S)$. Each node in the graph corresponds to a time stamp in the TS, while the edges between them represent their relationships. To determine the set of edges, Silva et al.[20] collected and presented three classes of approaches according to the underlying mapping concepts: the Visibility Graph Transformations, the Transition Networks and the Proximity Networks. Table 1 provides a comparative overview of these transformation methods. The first column lists all the variants that were utilised in this work. The second column specifies the node representation used in each algorithm, where $t$ denotes time points, while $q_{i}$, $\pi_{i}$ or $\bar{z}_{i}$ correspond to state based representations of the time points. The third column describes the edge type, such as Natural Visibility (NV) and Horizontal Visibility (HV) edges in Visibility Networks, Transition Probability (TP) edges indicating probability based connections, and Distance based Measure (DM) edges in Proximity Networks. The fourth and fifth columns indicate whether the generated networks have directed and weighted edges.
@@ 4.1 Visibility Networks
The visibility transformations from TS to complex networks are based on traditional visibility algorithms from computational geometry[21], associated with the natural ordering of the TS. In Table 1 we listed the algorithms that were utilised in this work.
The Natural Visibility Graph (NVG), originally introduced by Lacasa et al. [10], is based on the concept of visibility. In the NVG, each data point in the TS is treated as a vertical bar whose height corresponds to its value. Two points are connected if the line of sight between them is not obstructed by any intermediate data point. This means that there is a direct line of sight between the bars without any interception. Expressed mathematically: If any two points $(t_{i},s_{i})$ and $(t_{j},s_{j})$ have visibility, the following condition applies to every point $(t_{k},s_{k})$ in between:
@@
where: $t_{i}$, $t_{j}$ and $t_{k}$ are the time coordinates of the points, while $s_{i}$, $s_{j}$ and $s_{k}$ are the TS values at these specific time coordinates. The created edges $e$ are stored in an adjacency matrix $A^{NVG}$. The NVG algorithm is easy to implement and has quadratic complexity $O(S^{2})$[20] in terms of computational complexity. However, the NVG algorithm has two disadvantages, as it does not take into account the effect of uneven TS sampling and does not capture TS changes when the values are negative.
To overcome these disadvantages, a variation of NVG called the Weighted Dual Perspective Visibility Graph (WDPVG) was proposed by Zheng et al.[9]. According to the authors, the WDPVG method deals with uneven sampling in the time series by explicitly assigning weights to edges between nodes based on metrics such as Euclidean distance, tangent of the viewing angle, or actual time differences between connected points. Additionaly, WDPVG captures both peaks and troughs of the time series, by incorporating a dual perspective, which is achieved by combining the NVG with its reflected counterpart. To compute the reflected perspective of the NVG (RPNVG), the time series is inverted by mirroring its values on the time axis, resulting in a new series $S^{\prime}_{1\times N}=-S_{1\times N}$. A NVG is then constructed on this inverted series, resulting in an additional adjacency matrix $A_{RPNVG}$, which adds new edges $e_{i,j}$.
@@
Another version of the VG proposed by Luque et al.[11] is called the Horizontal Visibility Graph (HVG). Here, the visibility between two points is determined by a horizontal criterion. In a HVG, two nodes $v_{i}$ and $v_{j}$ are connected if all intermediate nodes $v_{k}$ satisfy the following condition:
@@
The Weighted [12] and Directed [22] Visibility Graphs are modifications of the core NVG or HVG, created by adding specific parameters to extend their functionality. By adding one of the two direction values: left to right (edge direction go from left to right according to the series temporal axis, which preserves the natural, temporal order of TSs) or top to bottom (edge directions go from top to bottom according to the value axis, focusing on magnitude differences between data points), we obtain a directed graph. Additionally, each edge in the graph going from one point to another can be associated with a feature vector or weight, denoted as $w_{ij}$. The ’weighted’ parameter can take one of the following values: Euclidean distance, squared Euclidean distance, horizontal or vertical distance and their absolute variants, slope, and absolute slope. The edges together with their weight values are stored in the adjacency matrix A.
@@ 4.2 Transition Networks
The computational complexity of the MTF increases quadratically with the length of the TS. Since the average length of the TS data in our benchmark datasets exceeds 3000, the application of the MTF becomes prohibitively expensive. Three main examples of this network type are Quantile Graphs[23], Ordinal Partition Transition Networks (OPTN) and Coarse-Grained Phase Space Graphs(CGPS).
@@ 4.3 Proximity Networks
The utilised type of the Procximity Network can be seen in the last row in Table 1. The proximity concept, on which these networks are based on, means that the measure of distance $d(s_{i},s_{j})$ or similarity (correlation) $r(s_{i},s_{j})$ is used to calculate the distance between the points in the TS incorporated in the multidimensional phase space. Nodes in this network represent the states of the TS, and edges are established based on the chosen distance or similarity metric. For example, the utilised k-NN network as shown in Table 1, has edges based on the measure of distance between the nodes. The ability to connect different nodes based on proximity measures allows capturing meaningful information about the topology of dynamical systems, enabling to identify different regimes or patterns along the series.[24]. The adjacency matrix for this type of network can, in general, be written as:
$A_{i,j}=\begin{cases}1&\text{if }d(s_{i},s_{j})\leq\epsilon\text{ or }r(s_{i},s_{j})\geq\alpha\text{ (binary matrix),}\\ w_{i,j}&\text{if proximity measure is used and weighting is applied},\\ 0&\text{otherwise}.\end{cases}$
Proximity networks include cycle networks[25][26], correlation[27] and recurrence networks[28].
@@ 5 Proposed segmentation model
In this section we describe our proposed model. As described in Section 3, our goal is a point-wise classification task, where a class must be assigned to each point in the TS. After transforming the TS into a graph representation, as described in Section 4, our point-wise classification task converts into a node classification task. In other words, each node in the graph corresponds to a time point, and we predict a label for each node. Formally, we define our prediction model as a function $\Phi$, which transforms the input data $G$ to the set of target classes $Y_{1\times N}$:

### v2

@@ 1 Introduction
In today’s digital era, time series (TS) are one of the most frequently encountered data types. Low-cost, high-resolution sensors have become a key part in mobile devices, industrial monitoring and healthcare, producing large amounts of TS data that capture the temporal patterns and essential statistical properties of various processes or phenomena [carvalho2019systematic]. With the rapid progress in automatic data acquisition and storage, we are now faced with the challenge of data that far exceeds our comprehension. TS analysis methods such as clustering, classification or regression, are therefore essential for effective and efficient processing of TS data [qaddoura2021multi, khan2018review, gers2000learning, aghabozorgi2015time]. Although these methods are crucial, they often benefit from first segmenting the TS into smaller, more manageable parts, through a process known as Time Series Segmentation (TSS). It first involves identifying critical time points dividing the TS into smaller segments and second label them based on activity. This segmentation step can significantly simplify subsequent tasks such as classification or anomaly detection and provide more targeted insights. For example, in the field of telecommunications [lebedev2023segmentation] the increasing volume of the network traffic data is a constant challenge. With TSS, we can segment network traffic, making it easier to detect patterns and respond swiftly to potential issues, such as peak usage periods, while optimizing bandwidth efficiency.
@@
In recent years, deep learning (DL) [perslev2019u, phan_8631195] architectures have been introduced as a TSS method, however they have not yet reached the same level of performance in TSS as in other fields such as computer vision. Many of these approaches rely on sliding-window techniques, where the TS is partitioned into evenly sized segments that are individually labeled before being recombined. While sliding-window methods yield satisfactory results, they limit the granularity of segmentation due to fixed window sizes and strides.
@@
A different, less mainstream approach to TS analysis, explores graphs as a viable TS representation [zheng2021visibility, lacasa2008time, luque2009horizontal, supriya2016weighted] where each value in the raw TS is represented as a node. Transforming TS into graphs not only preserves the temporal order of the data, but also reveal interactions and dependencies between non consecutive points, which is important for tasks such as TSS. This transformations enables the adoption of Graph Neural Networks (GNNs), which are particularly well-suited to learn complex structural features. GNNs can help identify the critical transition points that distinguish different segments, and are able to capture subtle temporal dependencies and variations that often precede significant changes in the TS. Such representations based on node density and message passing between node neighborhoods naturally eliminate the fixed window sizes and strides from DL.
@@
By considering the combination of TS graph representations and GNNs, and taking inspiration from the dense labeling approach in computer vision where every pixel is assigned a class label [shen2023survey], we can reformulate the TSS challenge as a node-level classification problem. In our graph representation of TS data, dense labeling entails assigning a specific label to each node, ensuring that even the most subtle transitions are detected and appropriately classified. In this work we aim to conduct a comprehensive and robust exploration of the different graph representations for TSS on the TSSB benchmark datasets in combination with GNNs. Based on the findings of this exploration, we propose a new approach that combines the Weighted Dual Perspective Visibility Graph (WDPVG) transformation with a GAT-based GNN architecture. The contributions of this paper are as follows:
@@ 2 Related work
Time series segmentation (TSS) is the task of partitioning time series into distinct, non-overlapping segments, each corresponding to an underlying state or pattern. Research in this area can be broadly categorized into three main categories depending on its application, as illustrated in Figure 1: Change Point Detection (CPD), which seeks to identify the specific time points of change; Windowed Segment Classification (WSC), which classifies fixed-length windows of the series; and Dense Labeling or Precise Segmentation (DLPS), which assigns a class label to every individual time step.
@@
As summarized in Table 1, these approaches have increased in complexity to address the limitations of earlier methods. The table lists key research works in the first column, followed by columns detailing the problem type addressed, the data type used, the number of datasets, and the core methodology employed.
Work Problem Type Data Type No. of Datasets Methodology
Arif et al. [arif2017change] CPD UTS 1 Statistical (CUSUM)
Ermshaus et al. [ermshaus2023clasp] CPD UTS 75 ML (k-NN Classifier)
De Ryck et al. [de2021change] CPD UTS 5 DL (Autoencoder)
Perslev et al. [perslev2019u] WSC MTS 7 DL (U-Time CNN)
Gaugel et al. [gaugel2023prectime] DLPS MTS 1 DL (CNN-LSTM)
This Work DLPS UTS (Graph) 59 DL (GNN)
Table 1: Summary of key related works.
@@
Early research predominantly focused on CPD. Initial approaches relied on statistical methods, such as those used by Arif et al. [arif2017change] to analyze climate data. Their methodology employed two statistical techniques: Cumulative SUM (CUSUM) to detect shifts in climate patterns, and bootstrapping to confirm the significance of those shifts. However, these statistical methods often struggle with subtle changes in autocorrelation or suffer from high false-alarm rates. This motivated the development of machine learning and deep learning solutions. For instance, De Ryck et al. [de2021change] employed a deep autoencoder to learn a stable, time-invariant representation of the time series. Change points are then identified by calculating a dissimilarity measure (e.g., the Euclidean distance) between consecutive windows’ feature representations, where notable increases in this measure indicate potential change points. In a different deep learning approach, Ermshaus et al. [ermshaus2023clasp] framed CPD as a self-supervised classification problem. Their method, ClaSP, uses a binary k-NN classifier to partition a time series recursively. For each potential split point, the classifier is trained to distinguish between the two resulting segments, and the split point yielding the highest accuracy is selected. ClaSP further automates this process by determining its own hyperparameters from the data, removing the need for manual tuning.
@@
However, a key limitation of CPD lies in its inability to classify the underlying nature of each segment. Consequently, TSS has evolved into a classification problem to overcome this gap. The most common strategy is WSC, in which the time series is split into fixed-length windows, each labeled independently. For instance, Perslev et al. [perslev2019u] employed a CNN-based U-Time architecture to segment sleep stages in EEG signals. A drawback, however, is that segmentation granularity depends on the chosen window size. Reducing the window to a single time stamp yields the most granular approach, referred to as DLPS. Gaugel et al. [gaugel2023prectime] introduced a deep learning model that merges sliding window and dense labeling techniques by employing CNNs for intra-window feature extraction, LSTM networks for inter-window contextual understanding, and a refinement module for dense labeling. Evaluated on multivariate time series with their model achieved a segmentation accuracy of approximately 96%.
This evolution reveals a clear and critical research gap, as shown in Table 1. While CPD for univariate time series (UTS) is well-explored, the more advanced WSC and DLPS methods have been overwhelmingly applied to multivariate time series (MTS) [perslev2019u, gaugel2023prectime]. The rich, multi-channel information in MTS data provides a strong feature basis for complex DL models. Univariate data lacks this explicit dimensionality, making it far more challenging to capture the complex temporal dependencies required for dense labeling.
To address this gap, our work formulates a novel approach to the DLPS of univariate data. We propose a dimensionality expansion technique that transforms the 1D time series into a graph representation. This allows us to leverage the powerful relational learning capabilities of Graph Neural Networks (GNNs). To the best of our knowledge, this work is the first to apply a GNN-based framework to the precise, dense labeling segmentation of univariate time series.
@@ 3 Problem formulation
TSS as a fundamental task in time series analysis involves partitioning a sequence into semantically meaningful segments. In this work, we formulate TSS as a point-wise classification problem, also referred to as DLPS, where each time step in a time series is assigned a categorical label indicating its segment membership. This formulation aligns well with graph node classification tasks [velickovic2018graph]. By representing the TS as a graph, where each time point corresponds to a graph node, the task of segmenting the series translates to assigning labels to these nodes.
@@ 4 Time series to graph transformations
In this section, we describe the main data preparation step, how to transform TS into graphs. We define the transformation $T()$ that transforms the input TS $S$ to its corresponding graph $G$: $G=T(S)$. Each node in the graph corresponds to a time stamp in the TS, while the edges between them represent their relationships. To determine the set of edges, Silva et al.[silva2021time] collected and presented three classes of approaches according to the underlying mapping concepts: the Visibility Graph Transformations, the Transition Networks and the Proximity Networks. Table 2 provides a comparative overview of these transformation methods. The first column lists all the variants that were utilised in this work. The second column specifies the node representation used in each algorithm, where $t$ denotes time points, while $q_{i}$, $\pi_{i}$ or $\bar{z}_{i}$ correspond to state based representations of the time points. The third column describes the edge type, such as Natural Visibility (NV) and Horizontal Visibility (HV) edges in Visibility Networks, Transition Probability (TP) edges indicating probability based connections, and Distance based Measure (DM) edges in Proximity Networks. The fourth and fifth columns indicate whether the generated networks have directed and weighted edges.
@@ 4.1 Visibility Networks
The visibility transformations from TS to complex networks are based on traditional visibility algorithms from computational geometry [ghosh2007visibility], associated with the natural ordering of the TS. In Table 2 we listed the algorithms that were utilised in this work.
The Natural Visibility Graph (NVG), originally introduced by Lacasa et al. [lacasa2008time], is based on the concept of visibility. In the NVG, each data point in the TS is treated as a vertical bar whose height corresponds to its value. Two points are connected if the line of sight between them is not obstructed by any intermediate data point. This means that there is a direct line of sight between the bars without any interception. Expressed mathematically: If any two points $(t_{i},s_{i})$ and $(t_{j},s_{j})$ have visibility, the following condition applies to every point $(t_{k},s_{k})$ in between:
@@
where: $t_{i}$, $t_{j}$ and $t_{k}$ are the time coordinates of the points, while $s_{i}$, $s_{j}$ and $s_{k}$ are the TS values at these specific time coordinates. The created edges $e$ are stored in an adjacency matrix $A^{NVG}$. The NVG algorithm is easy to implement and has quadratic complexity $O(S^{2})$[silva2021time] in terms of computational complexity. However, the NVG algorithm has two disadvantages, as it does not take into account the effect of uneven TS sampling and does not capture TS changes when the values are negative.
To overcome these disadvantages, a variation of NVG called the Weighted Dual Perspective Visibility Graph (WDPVG) was proposed by Zheng et al. [zheng2021visibility]. According to the authors, the WDPVG method deals with uneven sampling in the time series by explicitly assigning weights to edges between nodes based on metrics such as Euclidean distance, tangent of the viewing angle, or actual time differences between connected points. Additionaly, WDPVG captures both peaks and troughs of the time series, by incorporating a dual perspective, which is achieved by combining the NVG with its reflected counterpart. To compute the reflected perspective of the NVG (RPNVG), the time series is inverted by mirroring its values on the time axis, resulting in a new series $S^{\prime}_{1\times N}=-S_{1\times N}$. A NVG is then constructed on this inverted series, resulting in an additional adjacency matrix $A_{RPNVG}$, which adds new edges $e_{i,j}$.
@@
Another version of the VG proposed by Luque et al.[luque2009horizontal] is called the Horizontal Visibility Graph (HVG). Here, the visibility between two points is determined by a horizontal criterion. In a HVG, two nodes $v_{i}$ and $v_{j}$ are connected if all intermediate nodes $v_{k}$ satisfy the following condition:
@@
The Weighted [supriya2016weighted] and Directed [lacasa2012time] Visibility Graphs are modifications of the core NVG or HVG, created by adding specific parameters to extend their functionality. By adding one of the two direction values: left to right (edge direction go from left to right according to the series temporal axis, which preserves the natural, temporal order of TSs) or top to bottom (edge directions go from top to bottom according to the value axis, focusing on magnitude differences between data points), we obtain a directed graph. Additionally, each edge in the graph going from one point to another can be associated with a feature vector or weight, denoted as $w_{ij}$. The ’weighted’ parameter can take one of the following values: Euclidean distance, squared Euclidean distance, horizontal or vertical distance and their absolute variants, slope, and absolute slope. The edges together with their weight values are stored in the adjacency matrix A.
@@ 4.2 Transition Networks
The computational complexity of the MTF increases quadratically with the length of the TS. Since the average length of the TS data in our benchmark datasets exceeds 3000, the application of the MTF becomes prohibitively expensive. Three main examples of this network type are Quantile Graphs [de2015quantile], Ordinal Partition Transition Networks (OPTN) and Coarse-Grained Phase Space Graphs(CGPS).
@@ 4.3 Proximity Networks
The utilised type of the Procximity Network can be seen in the last row in Table 2. The proximity concept, on which these networks are based on, means that the measure of distance $d(s_{i},s_{j})$ or similarity (correlation) $r(s_{i},s_{j})$ is used to calculate the distance between the points in the TS incorporated in the multidimensional phase space. Nodes in this network represent the states of the TS, and edges are established based on the chosen distance or similarity metric. For example, the utilised k-NN network as shown in Table 2, has edges based on the measure of distance between the nodes. The ability to connect different nodes based on proximity measures allows capturing meaningful information about the topology of dynamical systems, enabling to identify different regimes or patterns along the series [donner2011recurrence]. The adjacency matrix for this type of network can, in general, be written as:
[…]
Proximity networks include cycle networks[zhang2006complex][zhang2006detecting], correlation[yang2008complex] and recurrence networks[marwan2008historical].
@@ 5 Proposed segmentation model
In this section, we describe our proposed model. As described in Section 3, our goal is a point-wise classification task, where a class must be assigned to each point in the TS. After transforming the TS into a graph representation, as described in Section 4, our point-wise classification task converts into a node classification task. In other words, each node in the graph corresponds to a time point, and we predict a label for each node. Formally, we define our prediction model as a function $\Phi$, which transforms the input data $G$ to the set of target classes $Y_{1\times N}$: